OpenStudy (anonymous):

Given the area of 1 side of the tissue box is x^2+6x+9 given by the formula A = Length times width (A=lw), what is the length of the tissue box?

OpenStudy (anonymous):

I don't even know... :0

OpenStudy (anonymous):

x^2+6x+9=lw ?

OpenStudy (anonymous):

This seems to be a factoring question, or so I think.

OpenStudy (anonymous):

yeah, that's what im doing

OpenStudy (anonymous):

What is the problem?

OpenStudy (anonymous):

the question I posted

OpenStudy (anonymous):

that's all I was given, haha

OpenStudy (anonymous):

I meant to say if you were having difficulties factoring or what?

OpenStudy (anonymous):

I don't even know how to start to solve this, how am I supposed to get the length??

OpenStudy (anonymous):

The only other hint I can tell you is that the length and the width are the same.

OpenStudy (anonymous):

I factored it already (x+3)(x+3) at least I think that's right...

OpenStudy (anonymous):

(x+3)(x+3)=LW

OpenStudy (anonymous):

Is that the answer that you're looking for?

OpenStudy (anonymous):

OH GOT IT, L=(x+3) silly me

OpenStudy (anonymous):

Good job